please help me solve the equation
3(4x+2)+5x=30-x

Found 2 solutions by edjones, stanbon:
Answer by edjones(8007)

3(4x+2)+5x=30-x
12x+6+5x=30-x
17x+6=30-x
18x=24
x=24/18=4/3
.
Ed

Answer by stanbon(75887)

3(4x+2)+5x=30-x
---
12x+6+5x = 30-x
17x+6 = 30-x
18x = 24
x = 4/3
================
Cheers,
Stan H.
=============